4. Solve for x, 0 SX < 360° a) 2 sinx + 3 = 0 b) 7cosx + 2 = 0 c) 2sinx - sinx = 0 d) 2cos2x - cosx - 1 = 0
Answer #1

a) 25in + 53 - 0 2 sim 2 = -53 sinx= -1 -2 X= sin(?) Sin O is negative in I & II quadrant so a would be [x = 300°) and x-240asiña - Sint=0 Sinx (2 sinx-1)=0 Sinx=0 2 Sinx=1 2= sin(0) SinX=lz x=sin (1) x = sin Es sin is quadrants positive in first a